# A new framework for nonlinear Kalman filters

### v6

@@ 2.2 The conventional framework and its problem
Theorem 1 suggests that using (9) to estimate the state covariance matrix will lead to overconfident covariance estimation when the estimated $S_{k}$ and $P_{xy,k}$ are not the same as their actual values. Again, this problem is unavoidable because the actual values of $S_{k}$ and $P_{xy,k}$ are related to the actual state, which is unknown. Note that this problem exists for all types of nonlinear KFs as long as (9) is used to update the state covariance estimation. We can further prove that the magnitude of such overconfidence is related to the accuracy of $S_{k|k-1}$ and $P_{xy,k|k-1}$. Namely, we have the following theorem.
Theorem 2.3.
Suppose that $P_{xy,k|k-1}=P_{xy}+t_{1}P,S_{k|k-1}=S_{k}+t_{2}S$, where $t_{1}\geq 0,t_{2}\geq 0$, $P$ is a random matrix with zero mean, and $S$ is a random symmetrical matrix with zero mean. Suppose that $P$ and $S$ are independent from each other, and $t_{2}S$ is very small. If we define the overconfidence in state covariance estimation as:
$\Delta p_{k|k}(t_{1},t_{2})=\operatorname{\mathbb{E}}[g(P_{xy,k},S_{k|k})-g(P_{xy,k|k-1},S_{k|k-1})]$ (15)
Then, for any $t_{a}\geq t_{b}$ and $t_{c}\geq t_{d}$, we have:
$\Delta p_{k|k}(t_{a},t_{c})\geq\Delta p_{k|k}(t_{b},t_{d})$ (16)
Proof 2.4.
Let $\Delta P=t_{1}P,\Delta S=t_{2}S$. (14) can be rewritten as:
$\Delta p_{k|k}(t_{1},t_{2})=\operatorname{\mathbb{E}}[g(P_{xy,k},S_{k|k})-g(P_{xy,k|k-1},S_{k|k-1})]$ (17)
$=$ $\operatorname{\mathbb{E}}[\text{tr}(2t_{2}^{2}(P_{xy,k}+t_{1}P)S_{k}^{-1}SS_{k}^{-1}SS_{k}^{-1}(P_{xy,k}+t_{1}P)^{T}$
$+2t_{1}^{2}PS_{k|k-1}^{-1}P^{T})]$
$=$ $2\operatorname{\mathbb{E}}[\text{tr}(t_{2}^{2}P_{xy,k}S_{k}^{-1}SS_{k}^{-1}SS_{k}^{-1}P_{xy,k}^{T}$
$+t_{1}^{2}t_{2}^{2}PS_{k}^{-1}SS_{k}^{-1}SS_{k}^{-1}P^{T}+t_{1}^{2}PS_{k|k-1}^{-1}P^{T})]$
Since $P_{xy,k}S_{k}^{-1}SS_{k}^{-1}SS_{k}^{-1}P_{xy,k}^{T}$, $PS_{k}^{-1}SS_{k}^{-1}SS_{k}^{-1}P^{T}$, and $PS_{k|k-1}^{-1}P^{T}$ are all positive semi-definite, their traces should all be greater or equal to zero. Therefore, (16) holds for any $t_{a}\geq t_{b}$ and $t_{c}\geq t_{d}$.
Theorem 2.3 suggests that increasing the noise in $S_{k|k-1}$ and $P_{xy,k|k-1}$ proportionally can make the state covariance estimation less accurate. In other words, the problem of the conventional framework is likely to be more severe when $S_{k}$ and $P_{xy}$ change more dramatically with respect to the states.
@@
In plain language, the problem of the conventional framework originates from the assumption that the selected Kalman gain is optimal. However, the Kalman gain, which was optimized based on the approximated measurement function near the predicted states, generally will not minimize the trace of the covariance matrix after the update. The reason is that the states will move to a new location after the update, and the previous approximations may be inaccurate. An example of such a problem is shown in Fig. 1, where we used an EKF to estimate a plane’s position. In the example, the state is the plane’s position along the x-axis, and the measurement function is the altitude of the ground beneath the plane, which is a nonlinear function of the state. Due to the inaccurate approximation of the measurement function around the predicted state (the dashed purple line, “Approx. 1”, in Fig. 1), the EKF overestimates the position of the plane, and the absolute state estimation error $|x_{k|k}-x_{k}|$ becomes higher after the “update” step. Meanwhile, the estimated standard deviation of the state, $\sigma_{k|k}$, becomes smaller after the “update” step. Such a contradiction suggests that the effectiveness of the Kalman gain is overestimated. This is visualized by the narrow yellow distribution in the “update” step in Fig. 1), which does not cover the true state.
@@ 2.3 The proposed new framework for nonlinear Kalman filters
Mathematically, “recalibrate” means to re-approximate the measurement functions around the updated states $x_{k|k}$ and re-estimate the matrices $P_{y,k},S_{k}$ and $P_{xy}$. We denote these re-approximated covariance matrices as $P_{y,k|k},S_{k|k}$, and $P_{xy,k|k}$ since they are all estimated based on the updated states. In the new framework, the state covariance matrix is updated as follows:
$P_{k|k}=$ $P_{k|k-1}+KS_{k|k}K^{T}-P_{xy,k|k}K^{T}-KP_{xy,k|k}^{T}$ (18)
$=$ $P_{k|k-1}+P_{xy,k|k-1}S_{k|k-1}^{-1}S_{k|k}S_{k|k-1}^{-1}P_{xy,k|k-1}^{T}$
$-P_{xy,k|k}S_{k|k-1}^{-1}P_{xy,k|k-1}^{T}-P_{xy,k|k-1}S_{k|k-1}^{-1}P_{xy,k|k}^{T}$
Since the new framework selects the same Kalman gain as the conventional framework, the actual value of the updated state covariance matrix $P_{k|k,ac}$ can still be calculated by (10). By comparing (18) and (10), we have the following theorem:
Theorem 2.5.
Assume that $\operatorname{\mathbb{E}}(P_{xy,k|k-1})=\operatorname{\mathbb{E}}(P_{xy,k|k})=P_{xy,k}$ and $\operatorname{\mathbb{E}}(S_{k|k-1})=\operatorname{\mathbb{E}}(S_{k|k})=S_{k}$. If $P_{xy,k|k-1}$, $P_{xy,k|k}$, $S_{k|k-1}$, and $S_{k|k}$ are all independent from each other, $P_{k|k}$ calculated in (18) is an unbiased estimation of $P_{k|k,ac}$. Namely, we have:
$\operatorname{\mathbb{E}}(P_{k|k})=P_{k|k,ac}$ (19)
Proof 2.6.
With the assumptions made in Theorem 2.5, we have:
$\operatorname{\mathbb{E}}(P_{k|k}-P_{k|k,ac})$ (20)
$=$ $\operatorname{\mathbb{E}}[P_{xy,k|k-1}S_{k|k-1}^{-1}(S_{k|k}-S_{k})S_{k|k-1}^{-1}P_{xy,k|k-1}^{T}]$
$-\operatorname{\mathbb{E}}[(P_{xy,k|k}-P_{xy,k})S_{k|k-1}^{-1}P_{xy,k|k-1}^{T}]$
$-\operatorname{\mathbb{E}}[P_{xy,k|k-1}S_{k|k-1}^{-1}(P_{xy,k|k}-P_{xy,k})^{T}]$
@@
Theorem 2.5 suggests that the proposed new framework can prevent the filter from being overconfident if the approximation errors of the measurement function at the predicted and updated states are independent of each other. This desirable feature comes from our replacement of (9) with (18), which highlights the fundamental idea of the new framework. That is, instead of reducing the updated state covariance matrix to the smallest possible value, we re-approximate the system after the state update to estimate the actual effect of the Kalman gain more accurately. In fact, in the new framework, the trace of the state covariance matrix may increase after the state update and recalibration if the system is strongly nonlinear and the predicted states have high variances, just like the example shown in Fig. 1. This phenomenon means that the update fails to make the estimation more accurate and should be withdrawn. To address this situation, we also add a “back out” step to the framework, which allows the state estimator to back out of the update when the update is unhelpful. The pseudo-code of the new framework is shown in Algorithm 1. Note that the new framework degrades to the conventional framework when the “recalibrate” and “back out” steps are replaced with (9).
@@
1:Input: Process noise covariance matrix $\boldsymbol{Q}_{k}$, Measurement noise covariance matrix $\boldsymbol{R}_{k}$, state transition function $\boldsymbol{f}(\boldsymbol{x},\boldsymbol{u})$, measurement function $\boldsymbol{h}{(\boldsymbol{x})}$, system inputs $\boldsymbol{u}_{k}$, measurements $\boldsymbol{z}_{k}$
@@
3:$\boldsymbol{\hat{x}}_{0|0}=\operatorname{\mathbb{E}}[\boldsymbol{x}_{0}]$
4:$\boldsymbol{P}_{0|0}=\operatorname{\mathbb{E}}[(\boldsymbol{\hat{x}}_{0|0}-\boldsymbol{x}_{0})(\boldsymbol{\hat{x}}_{0|0}-\boldsymbol{x}_{0})^{T}]$
@@
7:     Estimate $\boldsymbol{\hat{x}}_{k|k-1}$ and $\boldsymbol{P}_{k|k-1}$
8:    (based on $\boldsymbol{\hat{x}}_{k-1|k-1},\boldsymbol{P}_{k-1|k-1},\boldsymbol{f},\boldsymbol{u}_{k},$ and $\boldsymbol{Q}_{k}$)
@@
10:     Estimate $\boldsymbol{\hat{y}}_{k|k-1},\boldsymbol{P}_{xy,k|k-1}$ and $\boldsymbol{P}_{y,k|k-1}$
11:    (based on $\boldsymbol{\hat{x}}_{k|k-1},\boldsymbol{P}_{k|k-1},$ and $\boldsymbol{h}$)
12:     $\boldsymbol{S}_{k|k-1}=\boldsymbol{P}_{y,k|k-1}+\boldsymbol{R}_{k}$
13:     $\boldsymbol{K}=\boldsymbol{P}_{xy,k|k-1}\boldsymbol{S}_{k|k-1}^{-1}$
14:     $\boldsymbol{\hat{x}}_{k|k}=\boldsymbol{\hat{x}}_{k|k-1}+\boldsymbol{K}(\boldsymbol{\boldsymbol{z}_{k}-\hat{y}}_{k|k-1})$
@@
16:     Estimate $\boldsymbol{P}_{xy,k|k}$ and $\boldsymbol{P}_{y,k|k}$
17:    (based on $\boldsymbol{\hat{x}}_{k|k},\boldsymbol{P}_{k|k-1},$ and $\boldsymbol{h}$)
18:     $\boldsymbol{S}_{k|k}=\boldsymbol{P}_{y,k|k}+\boldsymbol{R}_{k}$
19:     $\boldsymbol{P}_{k|k}=\boldsymbol{P}_{k|k-1}+\boldsymbol{K}\boldsymbol{S}_{k|k}\boldsymbol{K}^{T}-\boldsymbol{P}_{xy,k|k}\boldsymbol{K}^{T}-\boldsymbol{K}\boldsymbol{P}_{xy,k|k}^{T}$
@@
21:     if $\text{tr}(\boldsymbol{P}_{k|k})>\text{tr}(\boldsymbol{P}_{k|k-1})$ then
22:         $\boldsymbol{\hat{x}}_{k|k}=\boldsymbol{\hat{x}}_{k|k-1}$
23:         $\boldsymbol{P}_{k|k}=\boldsymbol{P}_{k|k-1}$
@@
26:where, $\boldsymbol{\hat{x}}$ is the estimated states, $\boldsymbol{P}$ is the states’ covariance matrix, $\boldsymbol{\hat{y}}$ is the estimated measurements, $\boldsymbol{P}_{xy}$ is the covariance between the states and the measurements, $\boldsymbol{P}_{y}$ is the covariance of the estimated measurements, $\boldsymbol{S}$ is the innovation (or residual) covariance, $\boldsymbol{K}$ is the Kalman gain. The subscript $k|k-1$ represents the estimated value before the state update, and $k|k$ represents the estimated value after the state update.
It is worth mentioning that when the measurement functions are all linear, they would require no approximation. Or, to be more precise, the approximation of the measurement functions would become independent of the states. In this case, the estimations of $P_{y,k},S_{k}$, and $P_{xy,k}$ will not change before and after the recalibration, and our framework will be equivalent to the conventional framework.
To give readers an intuitive understanding of the effect of the “recalibrate” and “back out” steps, in Fig. 2, we showed an example of using different types of nonlinear KFs to do the state update given the measurement $y=x^{3}/3-x^{2}/8-x+1.5383=0$. The measurement function is represented by the thick gray curve in the figure. The predicted state is $x_{1|0}=0$, its variance is $1.5^{2}$, and the standard deviation of the measurement is $\sigma_{y}=0.01$. This state estimation problem is challenging because the derivative of the measurement function at the predicted state is negative, while the actual state is on the left of the predicted state. The state and covariance estimations made by different algorithms are summarized in different rows in the figure. Generally, the conventional framework always underestimates the actual state estimation error, which can be visualized by the narrow standard deviation estimations in Fig. 2 that cannot capture the true state. By contrast, the state covariance estimation made by the new framework is much more accurate. Additionally, the new framework is aware that the state estimation becomes less accurate after the update because the estimated variance increases. Therefore, after the “recalibrate” step, the “back out” step will be triggered, and the new framework will withdraw the update and use the predicted state and the predicted state variance as the final output.
@@
The new framework described above may remind the readers of the iterated Kalman filter (IKF), which also re-approximates the measurement function after calculating the Kalman gain [15, 16, 17]. However, IKF is fundamentally different from the new framework in three ways. Firstly, IKF redoes the approximation and recalculates the Kalman gain iteratively until convergence. However, such a convergence is not guaranteed, so the algorithm is less robust and often requires finetuning the step size of each iteration [18]. Even if the iteration converges, the convergence may take a very long time. By contrast, the new framework requires no iteration, and the re-approximation is only done once. Secondly, IKF was initially only used on EKF, and the deduction of iterated EKF (IEKF) requires an explicit approximation of the measurement functions (rows 9 and 10 in Algorithm 2), which is only possible for EKF and EKF2. Although a few attempts have been made to extend IKF to UKF and CKF [16, 17, 19, 20], a consensus on how to do this robustly has not been reached. Therefore, it is not easy to implement IKF on all types of nonlinear KFs while guaranteeing that the estimator can converge in most cases. Our framework, on the other hand, does not have this problem since there is no convergence issue. Thirdly, and most importantly, the basic idea behind IKF is still to find the optimal Kalman gain and not to estimate the covariance matrix accurately. Namely, IKF always updates the states and covariance matrix simultaneously using (9), resulting in overconfident state estimations. When the algorithm cannot converge properly, the estimation can be even worse because the estimated $P_{y,k},S_{k}$, and $P_{xy,k}$ can be less accurate, which can lead to a higher degree of overconfidence, according to Theorem 2.3. In general, in terms of making a more accurate estimation, finding the optimal Kalman gain is not as important as making the covariance estimation accurate, as we will show later in the simulation results.

### v10

@@ 2.2 The conventional framework and its problem
Theorem 1 suggests that using (10) to estimate the state covariance matrix tends to lead to overconfident covariance estimation when the estimated $S_{k}$ and $P_{xy,k}$ are not the same as their actual values. Admittedly, in practice, the errors of the approximated covariance matrices ${\color[rgb]{1,0,0}P_{y,k|k-1}},{\color[rgb]{1,0,0}S_{k|k-1}}$, and ${\color[rgb]{1,0,0}P_{xy,k|k-1}}$ are not as simple as zero-mean random matrices. However, Theorem 1 does give us a structured explanation as to why nonlinear KFs often underestimate the state covariance matrix. Note that an overconfident state covariance estimation can be very detrimental since it prevents the filter from correcting itself in subsequent iterations, thereby affecting the state estimation accuracy. Again, this problem is unavoidable because the actual values of $S_{k}$ and $P_{xy,k}$ are related to the actual state, which is unknown. Therefore, this problem exists for all types of nonlinear KFs as long as (10) is used to update the state covariance estimation.
@@
In plain language, the problem of the conventional framework originates from the assumption that the selected Kalman gain is optimal. However, the Kalman gain, which was optimized based on the approximated measurement function near the predicted states, generally will not minimize the trace of the covariance matrix after the update. The reason is that the states will move to a new location after the update, and the previous approximations may be inaccurate. An example of such a problem is shown in Fig. 1, where we used an EKF to estimate a plane’s position. In the example, the state is the plane’s position along the x-axis, and the measurement function is the altitude of the ground beneath the plane, which is a nonlinear function of the state. Due to the inaccurate approximation of the measurement function around the predicted state (the dashed purple line, “Approx. 1”, in Fig. 1), the EKF overestimates the position of the plane, and the absolute value of the state estimation error $|\hat{x}_{k|k}-x_{k}|$ becomes higher after the “update” step. Meanwhile, the estimated standard deviation of the state, $\sigma_{k|k}$, becomes smaller after the “update” step. Such a contradiction suggests that the effectiveness of the Kalman gain is overestimated. This is visualized by the narrow yellow distribution in the “update” step in Fig. 1, which does not cover the true state.
@@ 2.3 The proposed new framework for nonlinear Kalman filters
Mathematically, “recalibrate” means to re-approximate the measurement functions around the updated states $\hat{x}_{k|k}$ and re-estimate the matrices $P_{y,k},S_{k}$ and $P_{xy}$. We denote these re-approximated covariance matrices as ${\color[rgb]{1,0,0}P_{y,k|k}},{\color[rgb]{1,0,0}S_{k|k}}$, and ${\color[rgb]{1,0,0}P_{xy,k|k}}$ since they are all estimated based on the updated states. In the new framework, the state covariance matrix is updated by replacing $S_{k}$ with ${\color[rgb]{1,0,0}S_{k|k}}$ and $P_{xy,k}$ with ${\color[rgb]{1,0,0}P_{xy,k|k}}$ in (11):
${\color[rgb]{1,0,0}P_{k|k,new}}=P_{k|k-1}+{\color[rgb]{1,0,0}K_{est}}{\color[rgb]{1,0,0}S_{k|k}}{\color[rgb]{1,0,0}K_{est}^{T}}$ (18)
$\quad-{\color[rgb]{1,0,0}P_{xy,k|k}}{\color[rgb]{1,0,0}K_{est}^{T}}-{\color[rgb]{1,0,0}K_{est}}{\color[rgb]{1,0,0}P_{xy,k|k}^{T}}$
$=P_{k|k-1}+{\color[rgb]{1,0,0}P_{xy,k|k-1}}{\color[rgb]{1,0,0}S_{k|k-1}^{-1}}{\color[rgb]{1,0,0}S_{k|k}}{\color[rgb]{1,0,0}S_{k|k-1}^{-1}}{\color[rgb]{1,0,0}P_{xy,k|k-1}^{T}}$
$\quad-{\color[rgb]{1,0,0}P_{xy,k|k}}{\color[rgb]{1,0,0}S_{k|k-1}^{-1}}{\color[rgb]{1,0,0}P_{xy,k|k-1}^{T}}-{\color[rgb]{1,0,0}P_{xy,k|k-1}}{\color[rgb]{1,0,0}S_{k|k-1}^{-1}}{\color[rgb]{1,0,0}P_{xy,k|k}^{T}}$
Note that the proposed new framework selects the same Kalman gain and does the same state update as the conventional one; the only difference is that the state covariance update equation is now a direct approximation of (11), which is the actual state covariance regardless of the selection of the Kalman gain $K$. Intuitively, this can prevent the KF from overconfidence because the equation no longer relies on the problematic assumption that the Kalman gain is optimal. To explain this advantage more rigorously, we consider the newly approximated covariance matrices ${\color[rgb]{1,0,0}P_{y,k|k}},{\color[rgb]{1,0,0}S_{k|k}}$, and ${\color[rgb]{1,0,0}P_{xy,k|k}}$ as noisy versions of $P_{y,k},S_{k}$, and $P_{xy,k}$. Similar to our previous assumptions about ${\color[rgb]{1,0,0}P_{y,k|k-1}},{\color[rgb]{1,0,0}S_{k|k-1}}$, and ${\color[rgb]{1,0,0}P_{xy,k|k-1}}$, we assume that ${\color[rgb]{1,0,0}P_{xy,k|k}}$ is a random matrix, ${\color[rgb]{1,0,0}P_{y,k|k}}$ is a random positive semidefinite matrix, and ${\color[rgb]{1,0,0}S_{k|k}}$ is a random positive definite matrix. Additionally, we assume that the “noises” that arise in these approximations are unbiased, so these new random matrices should also satisfy:
$\begin{cases}\operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}P_{y,k|k}}]=P_{y,k}\\ \operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}S_{k|k}}]=S_{k}\\ \operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}P_{xy,k|k}}]=P_{xy,k}\end{cases}$ (19)
With these additional definitions and assumptions, we are able to give the following theorem about the advantage of the new framework:
Theorem 2.
Assume that $\operatorname{\mathbb{E}}({\color[rgb]{1,0,0}P_{xy,k|k-1}})=\operatorname{\mathbb{E}}({\color[rgb]{1,0,0}P_{xy,k|k}})=P_{xy,k}$ and $\operatorname{\mathbb{E}}({\color[rgb]{1,0,0}S_{k|k-1}})=\operatorname{\mathbb{E}}({\color[rgb]{1,0,0}S_{k|k}})=S_{k}$. If ${\color[rgb]{1,0,0}P_{xy,k|k-1}}$, ${\color[rgb]{1,0,0}P_{xy,k|k}}$, ${\color[rgb]{1,0,0}S_{k|k-1}}$, and ${\color[rgb]{1,0,0}S_{k|k}}$ are all independent from each other, ${\color[rgb]{1,0,0}P_{k|k,new}}$ calculated in (18) is an unbiased estimation of ${\color[rgb]{1,0,0}P_{k|k,ac}}$. Namely, we have:
$\operatorname{\mathbb{E}}({\color[rgb]{1,0,0}P_{k|k,new}}-{\color[rgb]{1,0,0}P_{k|k,ac}})=0$ (20)
{pf}
With the assumptions made in Theorem 2, we have:
$\operatorname{\mathbb{E}}({\color[rgb]{1,0,0}P_{k|k,new}}-{\color[rgb]{1,0,0}P_{k|k,ac}})$ (21)
$=$ $\operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}P_{xy,k|k-1}}{\color[rgb]{1,0,0}S_{k|k-1}^{-1}}({\color[rgb]{1,0,0}S_{k|k}}-S_{k}){\color[rgb]{1,0,0}S_{k|k-1}^{-1}}{\color[rgb]{1,0,0}P_{xy,k|k-1}^{T}}]$
$-\operatorname{\mathbb{E}}[({\color[rgb]{1,0,0}P_{xy,k|k}}-P_{xy,k}){\color[rgb]{1,0,0}S_{k|k-1}^{-1}}{\color[rgb]{1,0,0}P_{xy,k|k-1}^{T}}]$
$-\operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}P_{xy,k|k-1}}{\color[rgb]{1,0,0}S_{k|k-1}^{-1}}({\color[rgb]{1,0,0}P_{xy,k|k}}-P_{xy,k})^{T}]$
@@
The last equality equals zero because the terms in parentheses are zero, due to the unbiased (e.g. $\operatorname{\mathbb{E}}({\color[rgb]{1,0,0}S_{k|k}})=S_{k}$) and independence assumptions.
Theorem 2 suggests that the proposed new framework can generally prevent the filter from being overconfident if the approximation errors of the measurement function at the predicted and updated states are independent of each other. Admittedly, this assumption about the independence of the approximation errors is very hard to verify and may not hold in practice. Nonetheless, the theorem does explain why recalibrating the covariance matrices $P_{y,k},S_{k}$, and $P_{xy,k}$ after the state update is advantageous for accurate state covariance estimation. Specifically, the advantage comes from our replacement of (10) with (18), which highlights the fundamental idea of the new framework. That is, instead of reducing the updated state covariance matrix to the smallest possible value, we re-approximate the system after the state update to estimate the actual effect of the Kalman gain more accurately. In fact, in the new framework, the trace of the state covariance matrix may increase after the state update and recalibration if the system is strongly nonlinear and the predicted states have high variances, just like the example shown in Fig. 1. This phenomenon means that the update fails to make the estimation more accurate and should be withdrawn. To address this situation, we also add a “back out” step to the framework, which allows the state estimator to back out of the update when the update is unhelpful. The pseudo-code of the new framework is shown in Algorithm 1. Note that the new framework degrades to the conventional framework when the “recalibrate” and “back out” steps are replaced with (10).
@@
1:Input: Process noise covariance matrix $\bm{Q}_{k}$, Measurement noise covariance matrix $\bm{R}_{k}$, state transition function $\bm{f}(\bm{x},\bm{u})$, measurement function $\bm{h}{(\bm{x})}$, system inputs $\bm{u}_{k}$, measurements $\bm{z}_{k}$
@@
3:$\bm{\hat{x}}_{0|0}=\operatorname{\mathbb{E}}[\bm{x}_{0}]$
4:$\bm{P}_{0|0}=\operatorname{\mathbb{E}}[(\bm{\hat{x}}_{0|0}-\bm{x}_{0})(\bm{\hat{x}}_{0|0}-\bm{x}_{0})^{T}]$
@@
7:     Estimate $\bm{\hat{x}}_{k|k-1}$ and $\bm{P}_{k|k-1}$
8:    (based on $\bm{\hat{x}}_{k-1|k-1},\bm{P}_{k-1|k-1},\bm{f},\bm{u}_{k},$ and $\bm{Q}_{k}$)
@@
10:     Estimate $\bm{\hat{y}}_{k|k-1},\bm{P}_{xy,k|k-1}$ and $\bm{P}_{y,k|k-1}$
11:    (based on $\bm{\hat{x}}_{k|k-1},\bm{P}_{k|k-1},$ and $\bm{h}$)
12:     $\bm{S}_{k|k-1}=\bm{P}_{y,k|k-1}+\bm{R}_{k}$
13:     $\bm{K}=\bm{P}_{xy,k|k-1}\bm{S}_{k|k-1}^{-1}$
14:     $\bm{\hat{x}}_{k|k}=\bm{\hat{x}}_{k|k-1}+\bm{K}(\bm{z}_{k}-\bm{\hat{y}}_{k|k-1})$
@@
16:     Estimate $\bm{P}_{xy,k|k}$ and $\bm{P}_{y,k|k}$
17:    (based on $\bm{\hat{x}}_{k|k},\bm{P}_{k|k-1},$ and $\bm{h}$)
18:     $\bm{S}_{k|k}=\bm{P}_{y,k|k}+\bm{R}_{k}$
19:     $\bm{P}_{k|k}=\bm{P}_{k|k-1}+\bm{K}\bm{S}_{k|k}\bm{K}^{T}-\bm{P}_{xy,k|k}\bm{K}^{T}-\bm{K}\bm{P}_{xy,k|k}^{T}$
@@
21:     if $\text{tr}(\bm{P}_{k|k})>\text{tr}(\bm{P}_{k|k-1})$ then
22:         $\bm{\hat{x}}_{k|k}=\bm{\hat{x}}_{k|k-1}$
23:         $\bm{P}_{k|k}=\bm{P}_{k|k-1}$
@@
26:where, $\bm{\hat{x}}$ is the estimated states, $\bm{P}$ is the states’ covariance matrix, $\bm{\hat{y}}$ is the estimated measurements, $\bm{P}_{xy}$ is the covariance between the states and the measurements, $\bm{P}_{y}$ is the covariance of the estimated measurements, $\bm{S}$ is the innovation (or residual) covariance, $\bm{K}$ is the Kalman gain. The subscript $k|k-1$ represents the estimated value before the state update, and $k|k$ represents the estimated value after the state update.
It is worth mentioning that when the measurement functions are all linear, they would require no approximation. Or, to be more precise, the approximation of the measurement functions would become independent of the states. In this case, the approximations of $P_{y,k},S_{k}$, and $P_{xy,k}$ will not change before and after the recalibration, and our framework will be equivalent to the conventional framework.
To give readers an intuitive understanding of the effect of the “recalibrate” and “back out” steps, in Fig. 2, we showed an example of using different types of nonlinear KFs to do the state update given the measurement $y=x^{3}/3-x^{2}/8-x+1.5383=0$. The measurement function is represented by the thick gray curve in the figure. The predicted state is $x_{1|0}=0$, its variance is $1.5^{2}$, and the standard deviation of the measurement is $\sigma_{y}=0.01$. This state estimation problem is challenging because the derivative of the measurement function at the predicted state is negative, while the actual state is to the left of the predicted state. The state and covariance estimations made by different algorithms shown in different rows in the figure. Generally, the conventional framework always underestimates the actual state estimation error, which can be visualized by the narrow standard deviation estimations in Fig. 2 that cannot capture the true state. By contrast, the state covariance estimation made by the new framework is much more accurate. Additionally, the new framework is aware that the state estimation becomes less accurate after the update because the estimated variance increases. Therefore, after the “recalibrate” step, the “back out” step will be triggered, and the new framework will withdraw the update and use the predicted state and the predicted state variance as the final output.
@@
2.4 Relationship to Iterated Kalman Filter
The new framework described above may remind the readers of the iterated Kalman filter (IKF), which also re-approximates the measurement function after calculating the Kalman gain [19, 20, 21]. However, IKF is fundamentally different from the new framework in three ways. Firstly, IKF redoes the approximation and recalculates the Kalman gain iteratively until convergence. However, such a convergence is not guaranteed, so the algorithm is less robust and often requires finetuning the step size of each iteration [22]. Even if the iteration converges, the convergence may take a very long time. By contrast, the new framework requires no iteration, and the re-approximation is only done once. Secondly, IKF was initially only used on EKF, and the deduction of iterated EKF (IEKF) requires an explicit approximation of the measurement functions (rows 9 and 10 in Algorithm 2), which is only possible for EKF and EKF2. Although a few attempts have been made to extend IKF to UKF and CKF [20, 21, 23, 24], a consensus on how to do this robustly has not been reached. Therefore, it is not easy to implement IKF on all types of nonlinear KFs while guaranteeing that the estimator can converge in most cases. Our framework, on the other hand, does not have this problem since there is no convergence issue. Thirdly, and most importantly, the basic idea behind IKF is still to find the optimal Kalman gain and not to estimate the covariance matrix accurately. Namely, IKF always updates the states and covariance matrix simultaneously using (10), resulting in overconfident state estimations. When the algorithm cannot converge properly, the estimation can be even worse because the estimated $P_{y,k},S_{k}$, and $P_{xy,k}$ can be less accurate, which can lead to a higher degree of overconfidence. In general, in terms of making a more accurate estimation, finding the optimal Kalman gain is not as important as making the covariance estimation accurate, as we will show later in the simulation results.